Of the drivers who stop at a gas station, 91% purchase gasoline, and 6% purchase both gasoline and oil. A total of 9% purchase oil.
(a) What is the probability that a driver purchases gasoline, given that he or she purchases oil? Round your answer to decimal places.
(b) What is the probability that a driver purchases oil, given that he or she purchases gasoline? Round your answer to decimal places.

Answer by edjones(7569)

n=intersection, G=gas, O=oil
.
a)P(G|O)=P(GnO)/P(O)=.06/.09=2/3
.
b)P(O|G)=.06/.91=.0659
.
Ed